OpenStudy (anonymous):

Evaluate the expression pi/6 over 1/4 i really dont understand

OpenStudy (anonymous):

Do you mean $$ \frac{\frac{\pi}{6}}{4} $$ ?

OpenStudy (anonymous):

theres a 1 over the 4

OpenStudy (dangerousjesse):

Basically what they're saying is \[\frac{ \frac{ \pi }{ 6 } }{ \frac{ 1 }{ 4 } }\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorry, I read that wrong

OpenStudy (mathmale):

Look at pi/6 (itself a fraction) as the numerator of a new fraction; look at 1/4 as the denominator of this new fraction...exactly as DangerousJesse has shown you.

OpenStudy (mathmale):

What is the rule for dividing one fraction by another (as you must do here)?

OpenStudy (anonymous):

flip?

OpenStudy (mathmale):

Take the denominator, flip it, and then multiply the result by the numerator. Here, take the denominator, 1/4, flip it over (to get 4/1), and multiply that by the numerator pi/6.

OpenStudy (anonymous):

\[(\frac{ \pi }{ 6 }) \div (\frac{ 1 }{ 4 }) = (\frac{ \pi }{ 6 }) \times (\frac{ 4 }{ 1 })\]

OpenStudy (mathmale):

@CoolRanchDoritos : Go for it!

OpenStudy (anonymous):

thank you vary much

OpenStudy (anonymous):

ok:) ur welcome